OpenStudy (anonymous):

what is 4=16^x / 4^x ?????????

OpenStudy (blurbendy):

\[4 = \frac{ 16^x }{ 4^x }\] It looks like we can simplify the right side pretty nicely. Any ideas?

OpenStudy (jdoe0001):

\(\bf 4 = \cfrac{16^x}{4^x} \implies 4\times 4^x = 16^x\\ 4^{x+1} = 16^x\)

OpenStudy (jdoe0001):

from there try to use the log cancellation rule of => \(\bf\huge log_aa^x = x\)

OpenStudy (jdoe0001):

confused still @vramirez83 ?

OpenStudy (blurbendy):

I think this problem is made easier by noticing that 16^x / 4^x = 4^x So, we have 4^x = 4 instead of taking a log here, what number (x), gives us 4 when we take 4 to the x power?

OpenStudy (anonymous):

no i got it thank you (: